OpenStudy (anonymous):

Members of a school social service group hold a spaghetti dinner to raise money for orphan children. They must sell tickets at a price that will earn them more money than the cost of food. The formula for this scenario is t = n (p − c) where t = total profit earned from the event c = cost of food per plate n = number of tickets sold p = price charged for each dinner Rewrite the equation to determine the number of tickets they need to sell.

OpenStudy (anonymous):

So in other words, you need to rearrange \[t = n(p-c)\] So n is the subject. Does that make it easier?

OpenStudy (anonymous):

Price - cost x number of tickets = answer

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

help

OpenStudy (anonymous):

I'll start it off for you. \[t = n(p-c)\] Expand the brackets \[t = np - nc\] Now rearrange for n

OpenStudy (anonymous):

laura shes not gonna give you an answer